OpenStudy (anonymous):

Determine the sum of the infinite geometric series 1/7+(1/7)^2+(1/7)^3+...

OpenStudy (slaaibak):

a/1-r a=1/7 r=1/7 (1/7)/(6/7) = 1/6

OpenStudy (anonymous):

Thank you!!